# Uniform Field in Microwave Cavities Through the Use of Effective Magnetic Walls

### v1

@@ Abstract
Figure 1: Schematic of the proposed modification to the cavity’s boundary condition. The WM resonator consists of metallic wires with radius $r$ and period $a$. In the regular configuration, the distance between the outermost wires and the PEC walls is $a/2$. Adjusting this distance to $a/2+\lambda/4$ transforms the boundary, making it behave as an effective perfect magnetic conductor wall.
@@
Wire media (WM), introduced by Brown Brown (1960) and Rotman Rotman (1962) and later advanced by Pendry and Belov et al. Pendry et al. (1998); Belov et al. (2003), are artificial plasma materials with tunable plasma frequencies determined by their geometric parameters. These dense arrays of metallic wires embedded in a dielectric host have been applied in diverse fields, ranging from microwave antennas Forati, Hanson, and F. Sievenpiper (2015) to near-field thermophotovoltaic systems Simovski et al. (2013), and systems of radiative cooling Simovski et al. (2015) (see Ref. Simovski et al., 2012 for a detailed review).
@@
Research on WM open resonators has primarily focused on utilizing topological transitions in isofrequency contours to enhance the Purcell factor Mirmoosa, Kosulnikov, and Simovski (2016a, b). Recently, WM cavities enclosed by metallic walls and placed in a DC magnetic field—known as plasma haloscopes Lawson et al. (2019)—have emerged as a promising technique for detecting axions, a dark matter candidate Millar et al. (2023). These cavities can also be used to search for dark photons, another dark matter candidate, even in the absence of a magnetic field Gelmini et al. (2020). Our previous work demonstrated that WM operating in the epsilon-near-zero (ENZ) regime significantly extends the wavelength of light, achieving low-frequency resonance and high field uniformity Balafendiev et al. (2022).
@@
Field homogeneity in microwave resonators has been studied in the context of electron paramagnetic resonance (EPR), where axially uniform TE modes are desirable Hyde, Sidabras, and Mett (2019). A common method involves a central metallic waveguide operating at its cutoff frequency, flanked by end regions with optical lengths set to a quarter wavelength Hyde, Sidabras, and Mett (2019). Proposed configurations for the end regions include dielectric rods matching the central cross-section Mett, Froncisz, and Hyde (2001), enlarged cross-sections Anderson, Mett, and Hyde (2002), and conducting rod insertions Hyde, Mett, and Anderson (2002).
@@
For axion haloscopes, field uniformity is as beneficial as it is for EPR devices. Key characteristics that enhance performance include the strength of the DC magnetic field, cavity volume, quality factor, and form factor, all of which influence the power from photon-axion conversion in resonance Krauss et al. (1985). The form factor for a mode in a plasma haloscope, characterized by an electric field $\mathbf{E}_{lmn}$, is defined as
@@
where $V$ is the cavity volume, $\epsilon$ is the relative permittivity within the cavity, and $\mathbf{B}_{0}$ is the external DC magnetic field, assumed spatially constant and aligned with the wires.
To achieve optimal field homogeneity and form factor in plasma haloscopes, researchers have focused on the fundamental TM cavity mode Wuensch et al. (1989); Bradley et al. (2003); Asztalos et al. (2001); Stern et al. (2019). In WM resonators, the fundamental TM mode produces a uniform axial field and a sinusoidal transverse pattern due to electric wall boundary conditions and axial polarization Balafendiev et al. (2022). Wedge-shaped haloscopes use quarter-wavelength corrugations in the walls to achieve axial field uniformity Dyson et al. (2024); Kuo (2020). However, researchers have yet to explore improving field homogeneity in the transverse direction of WM resonators.
In this work we demonstrate that placing the cavity walls with a quarter-wavelength gap from the WM interface transforms the interface from a perfect electric conductor to a perfect magnetic conductor. Figure 1 illustrates this concept. This transformation is similar to converting a shortened transmission line into an open one by shifting the termination by an odd number of quarter-wavelengths. By modifying the boundary conditions, we create a cavity with a $\mathrm{TM}_{000}$ mode instead of the regular $\mathrm{TM}_{110}$ mode, significantly improving the form factor and lowering the fundamental resonance frequency to the plasma frequency.
Figure 2: One-dimensional schematic of a resonator consisting of a wire medium (region 1) with thickness $d$, enclosed by an air gap (region 2) of thickness $g$.
To evaluate the impact of this air gap on modal field uniformity, we develop a one-dimensional analytical model that focuses on TM modes, which interact most strongly with the WM. We model the system as two regions, as shown in Fig. 2. Region 1 contains the lossless WM (PEC wires) characterized by a relative permittivity Simovski et al. (2012)
@@
where $k_{x1}=\sqrt{\epsilon}k_{0}$, and $k_{x2}=k_{0}$, assuming no field dependence on $x$ and $z$.
We derive the surface impedance for the WM-air interface by evaluating the field solutions from Eqs. (3) and (4) at $x=\tfrac{d}{2}^{+}$, just outside the WM region, yielding the following expression:
@@
This impedance vanishes for $g=0$, corresponding to a perfect electric conductor, and approaches infinity as $g$ tends to $\lambda/4$, matching the wave impedance of a perfect magnetic conductor.
Building on the derived surface impedance, we now investigate the TM modes of a 2D WM resonator using the 1D model. Assuming no field variation along the $z$-axis for the fundamental mode ($k_{z}=0$) and a square resonator cross-section ($k_{x}=k_{y}=k_{t1}$), the electromagnetic field within the effective WM is given by
@@
In the effective WM region, defined by $|x|,\>|y|\leq\tfrac{d}{2}$, we define the wave number as
@@
while in the air gap region, for instance where $x>\tfrac{d}{2},|y|\leq\tfrac{d}{2}$, we define the wave number as
@@
This model assumes the WM resonator has no corners, which may differ from a real resonator. However, we can use the one-dimensional solution to estimate the two-dimensional behavior, similar to Marcatili’s method for dielectric rectangular waveguides Marcatili (1969). This approximation holds when the field is mainly concentrated in the central region.
@@
This equation gives the resonance frequencies of the TM modes. Notably, using Eq. (9), we can determine the resonance frequency of the fundamental TM mode. For the fundamental mode, when $g=\lambda/4$, the second term in Eq. (10) vanishes, implying that $k_{t1}=0$. Consequently, $\epsilon=0$ and $k_{0}=k_{p}$, leading to a constant electric field according to Eq. (6).
It is worth mentioning that Eq. (10) resembles the dispersion equation derived in Ref. Mett, Froncisz, and Hyde, 2001. However, while Ref. Mett, Froncisz, and Hyde, 2001 achieves field homogeneity in only one direction using a central metallic waveguide, our approach offers the advantage of ensuring field homogeneity in both transverse directions by utilizing a WM in the central region of the cavity.
@@
Figure 3: Electric field distribution in a WM resonator: (a) Regular configuration with no air gap ($g=0$) and (b) configuration with an air gap of $g=\lambda_{p}/4$. Introducing the air gap improves the uniformity of the electric field distribution.
@@
Figure 4: WM resonator electric field profile for different air gap thicknesses, $g$. The WM extends from $x=-\tfrac{d}{2}$ to $x=\tfrac{d}{2}$. A uniform electric field within the WM region is achieved at $g=\lambda_{p}/4$.
@@
Figure 5: Characteristics of the fundamental TM mode in a WM resonator: (a) Resonance frequency and (b) form factor as functions of air gap thickness, $g$. At $g=\lambda_{p}/4$, a form factor of approximately 0.92 is achieved for the effective WM (theory and simulations), while a value of 0.89 is observed for the physical WM (simulations). Point P indicates the air gap thickness applied in the experimental setup.
@@
We validated the applicability of our theoretical model using wave simulations with the COMSOL Multiphysics eigenmode solver. Both our analytical calculations and simulations modeled the cavity as an array of 10$\times$10 infinitely long wires, with a wire radius of $r_{0}$ = 0.5 mm and inter-wire spacing $a$ = 10 mm. To determine the required plasma frequency for the theoretical model, we applied a generalized quasi-static approach Kumar et al. (2012).
@@
Figure 3 shows the normalized electric field distribution for the fundamental TM mode over the transverse ($xy$) plane of the cavity for two cases: $g=0$ and $g=\lambda_{p}/4$. Both our numerical and analytical results confirm that introducing a $\lambda_{p}/4$ air gap homogenizes the electric field distribution in a finite-cross section sample of the WM.
@@
The electric field profile along the $x$-axis for the fundamental TM mode of a WM resonator varies with the gap width, $g$, as shown in Fig. 4. The field homogeneity improves as $g$ increases, reaching a maximum at $g=\lambda_{p}/4$. Introducing a quarter-wavelength air gap between the effective wire medium and the PEC walls significantly flattens the electric field profile, confining the field decay primarily to the newly added $\lambda_{p}/4$ region. This coincides with a pronounced magnetic field enhancement within the same gap, offering potential for readout applications. When the gap exceeds $\lambda_{p}/4$, the field concentrates near the resonator walls, and the underlying assumption of field confinement within the wire medium becomes not adequate.
@@
Figure 5 illustrates the resonant frequency (Fig. 5a) and form factor (Fig. 5b) of the fundamental mode in such WM resonator as a function of the air gap, $g$. The resonant frequency remains nearly constant for $40\leq g\leq\lambda_{p}/$ 4, aligning with the plasma frequency at $g=\lambda_{p}/4$. Conversely, the form factor progressively increases, reaching a maximum at an air gap of approximately $g=\lambda_{p}/4$. While the resonant frequency exhibits excellent agreement between analytical and numerical results, coming from both a physical WM and an effective WM, within this range, a slight discrepancy arises in the form factor. This error results from the evident difference between the effective and physical (real) WM. In the effective-medium model there are no wires and the field does not vanish inside this medium, whereas in the physical WM it vanishes within every wire – see the dark spots in the bottom panel of Fig. 3b.
Numerical calculations of the mode field for an effective WM show a maximum form factor of 0.92 at 8.9 GHz with $g=\lambda_{p}/4=8.4$ mm, consistent with the analytical model. However, the simulation of the physical WM corresponding to the bottom panel of Fig. 3b shows the form factor 0.89. Despite this, both numerical and analytical results agree on the air gap value for the maximum form factor. For smaller gaps the agreement between the analytical and numerical results corresponding to the physical WM is good, and is excellent for the numerical simulations of the effective WM sample surrounded by the PEC walls. Thus, these results fully justify the model of an effective WM sample for $g\leq\lambda_{p}/4$. For $g>\lambda_{p}/4$, the analytical model diverges from numerical simulations due to field concentration in the air gap, which invalidates the assumption of field confinement within the WM. Nonetheless, numerical results for the effective WM and the array of separate wires remain consistent.
@@
The maximal form factor corresponds to the perfect magnetic wall effectively located at the WM-air interface. The PEC wall shifts the phase of a reflected wave by $\pi$, the air gap grants an extra $\pi$, so the phase of the reflected wave does not change. Therefore, the mode inside the WM sample is the standing wave shaped by four perfect magnetic boundaries surrounding the sample. Note that this effective boundary condition can be achieved through various methods that introduce an extra phase shift of $\pi$ between incident and reflected waves, as detailed in Mett, Froncisz, and Hyde (2001); Anderson, Mett, and Hyde (2002); Hyde, Mett, and Anderson (2002). This flexibility may facilitate the implementation of TM000 modes in tunable plasma haloscope designs.
@@
Figure 6: Prototype WM resonators with different wall spacings, consisting of 10 x 10 arrays of copper wires with a period of $a=1$ cm. The first resonator is regular with a distance of $a/2$ between the edge wires and the walls, and the second is $\lambda/4.2$-shifted, with spacing $a/2+\lambda/4.2$. (a) Unfinished state of the regular resonator with copper wires (radius 0.5 mm) inserted into the FR-4 PCB walls. (b) Finished prototype being scanned through pre-drilled holes by a movable antenna.
@@
To validate the claimed effect, we fabricated two prototypes. The first prototype represents the regular (conventional) configuration where the air gap between the array of wires and the cavity walls is $a/2$ Balafendiev et al. (2022). The second prototype incorporates a configuration we specially optimized for the cubic resonator to enhance field homogeneity and to avoid situations where the air gap exceeds $\lambda_{p}/4$, as this would lead the field to localize mainly near the cavity walls (see Fig. 4). In this 3D case, the designed gap is slightly smaller than the quarter-wavelength. So, for the optimized resonator, the edge wires of the array are distanced from the walls by ($a/2+\lambda_{p}/4.2$). For the 2D resonator, this configuration would correspond to point P in Fig. 5. Below we call our optimized cubic resonator the $\lambda_{p}/4.2$-shifted one.
@@
We fabricated the prototype enclosures using single-sided metal FR-4 PCBs of $1$ mm thickness. We arranged these PCBs into cubes with dimensions of 10 cm x 10 cm x 10 cm for the regular resonator and 11.6 cm x 11.6 cm x 11.6 cm for the $\lambda/4.2$-shifted resonator. To complete the resonators, we inserted and solded one hundred copper wires, each with a radius of 0.5 mm, into the PCB cubes with a periodic spacing of $a=1$ cm, as shown in Fig. 6a.
@@
Figure 7: Experimentally measured $\mathrm{S}_{21}$ parameter for WM resonators: (a) Regular resonator and (b) $\lambda_{p}/4.2$-shifted resonator. Measurements were performed by placing the movable monopole antenna in the central hole. The resonance frequency of the fundamental TM mode in the regular resonator ($f_{110}$) is $9.11$ GHz (experiment) and $9.10$ GHz (simulation), while in the $\lambda_{p}/4.2$-shifted resonator ($f_{000}$) it is $8.94$ GHz (experiment and simulation).
We measured the S-parameters using a Rohde &\&& Schwarz ZVB20 vector network analyzer (VNA) across a frequency range of 8.75-9.75 GHz. The fixed antenna on the bottom plate of each prototype was connected to one VNA port, while the movable antenna on the top plate was connected to the other port. Fig. 7 displays the measured $\mathrm{S_{21}}$ when the movable antenna was placed in the central hole of each prototype. We observe that the spectrum of the $\lambda_{p}/4.2$-shifted resonator is red-shifted relative to that of the regular resonator, as predicted by theory for the fundamental mode. The amplitude levels and spectrum shape change slightly, suggesting that the quality factor is not significantly affected by the introduction of a $\lambda_{p}/4.2$ air gap.
We conducted full-wave simulations of cubic resonators, matching the experimental prototypes, using CST Studio. Copper was modeled with a conductivity of $\sigma=5.7\times 10^{7}\;\mathrm{S/m}$, while the FR-4 layers of the PCBs were represented as lossy dielectrics with a dielectric constant of $\varepsilon=4.3$ and a loss tangent of $\tan{\delta}=0.025$. We measured the $\mathrm{S_{21}}$ parameter and positioned a field monitor at the first resonance frequency for each resonator.
We characterized the operation of the two resonators under comparison by examining the first peaks of the $\mathrm{S_{21}}$-parameter spectra. These peaks correspond to the $\mathrm{TM_{110}}$ mode for the regular resonator and the quasi-$\mathrm{TM_{000}}$ mode for the $\lambda_{p}/4.2$-shifted resonator, as shown in Fig. 7. Fig. 8 presents the measured electric field distributions for both prototypes at their corresponding frequencies ($f_{000}$ for the $\lambda_{p}/4.2$-shifted resonator and $f_{110}$ for the regular one). We observe a significant improvement in field homogeneity for the $\lambda_{p}/4.2$-shifted resonator. This observation is further supported by the measured form factors: $0.77$ (experiment), $0.80$ (simulation) for the $\lambda_{p}/4.2$-shifted resonator compared to $0.61$ (experiment), $0.64$ (simulation) for the regular resonator.

### v3

@@ I Introduction
Figure 1: Schematic representation of the proposed modification to the cavity’s boundary condition. A wire medium (WM) resonator consists on an arrange of metallic wires with radius $r$, distributed with a period $a$. In a regular configuration, the distance between the last set of wires and the cavity walls, made of perfect electric conductor (PEC), is $a/2$. By adjusting this distance to $a/2+\lambda/4$, the boundary enclosing the WM behaves effectively as a perfect magnetic conductor (PMC) wall.
@@
Wire media (WM), initially proposed by Brown [1] and Rotman [2], and later studied by Pendry and Belov et al. in works [3, 4], have garnered significant attention for their distinctive properties as an artificial plasma material. These structures, comprising a dense array of parallel metallic wires embedded in a dielectric host, exhibit a tunable plasma frequency that can be adjusted by modifying their geometrical parameters. Beyond theoretical exploration, WM-based devices have found practical applications in diverse fields, ranging from microwave antennas [5] to near-field thermophotovoltaic systems [6] and systems of radiative cooling [7]. A detailed review about WM applications can be found in Ref. [8].
@@
Previous investigations into WM open resonators have primarily focused on exploiting topological transitions in their isofrequency contours to achieve a notable Purcell factor [9, 10]. More recently, cavities consisting in WM enclosed by metallic walls and placed in a DC magnetic field, often referred to as plasma haloscopes [11], have emerged as a promising technique for the detection of axions, a possible candidate for dark matter [12]. One can also look for dark photons, another promising dark matter candidate even if the magnetic field is absent [13]. Our earlier studies of WM-filled cavities demonstrated that WM operating in the epsilon-near-zero (ENZ) regime significantly extends the wavelength of light within the medium, resulting in a low-frequency fundamental resonance and a high degree of field uniformity [14].
@@
The improvement of field homogeneity in microwave resonators have been studied in the context of electron paramagnetic resonance (EPR), where axially uniform field distribution for a TE mode is a desirable feature [15]. A common approach to higher uniformity of the field pattern in hollow metal cavities involves a design with a central region surrounded by two end regions. The central region functions as a metallic waveguide operating at its cutoff frequency, while the optical length of the end regions is set to a quarter of a wavelength [15]. Various configurations for end regions in hollow metal cavities have been proposed, including dielectric rods with the same cross-section as the central region [16], modified cross sections larger than that of the central region [17], and conducting rods insertions [18].
@@
As to the axion haloscopes, the field uniformity is just as beneficial for them as for EPR devices. Several key characteristics should be engineered to enhance their performance: the strength of the DC magnetic field, cavity volume, quality factor, and form factor are critical factors influencing the power from photon-axion conversion in resonance [19]. The form factor for a mode in a plasma haloscope, characterized by an electric field $\mathbf{E}_{lmn}$, is defined as
@@
where $V$ is the cavity volume, $\epsilon$ is the relative permittivity within the cavity, and $\mathbf{B}_{0}$ is the external DC magnetic field, which we will take to be spatially constant and aligned with the wires.
To ensure the highest field homogeneity inside plasma haloscopes, and thus the optimal form factor, the focus has been on the study of the fundamental TM cavity mode [20, 21, 22, 23]. The fundamental TM mode in a WM resonator, with WM as the cavity content, exhibits a uniform axial field profile and a sinusoidal transverse pattern due to the electric wall boundary conditions and the axial polarization of the electric field [14]. In wedge-shaped haloscopes, the quarter-wavelength long corrugations of the walls were utilized to achieve the field uniformity in the axial direction [24, 25]. Meanwhile, improving field homogeneity in WM resonators along the transversal section remains an unexplored area.
In this work we show that placing the walls with the quarter-wavelength gap from the WM interface, we effectively transform this interface from the perfect electric conductor (PEC) into a perfect magnetic conductor (PMC). Figure 1 illustrates this concept. This transformation is similar to the conversion of a shortened transmission line into an open one by shifting the termination by an odd number of quarter-wavelengths. By modifying the boundary conditions, we create a cavity whose fundamental mode is $\mathrm{TM}_{000}$ rather than the $\mathrm{TM}_{110}$ mode of the regular cavity. This transition significantly improves the form factor and shifts the fundamental resonance frequency down to the plasma frequency.
To investigate the enhancement of form factor in wire medium (WM) resonators, we propose an analytical model taking into account the boundary conditions at the WM interfaces separated from the cavity walls by arbitrary air gaps. Using this model, we estimate the resonance frequency and form factor of square transverse section WM cavities. We compare the analytical results with numerical simulations and present experimental and full-wave simulation data for the spectra and form factors of two WM resonator prototypes fabricated using printed circuit boards (PCBs). Our findings demonstrate a significant increase in the form factor when introducing an air gap with an optimal thickness, close to a quarter of the resonant wavelength, between the WM and the cavity walls.
II Analytics
II.1 Effective boundary condition
Figure 2: One-dimensional schematic of a resonator composed of a wire medium (region 1) with thickness $d$ surrounded by an air gap (region 2) with thickness $g$.
To assess the impact of the air gap between the wall and the WM on the modal field uniformity, we develop a one-dimensional analytical model. We focus on the TM modes because the TE modes weakly interact with the WM. We divide the system into two regions, as depicted in Fig. 2. Region 1 contains the lossless (PEC wires) WM characterized by a relative permittivity [8]
@@ II.1 Effective boundary condition
where $k_{x1}=\sqrt{\epsilon}k_{0}$, and $k_{x2}=k_{0}$, assuming no field dependence on $x$ and $z$. We evaluate the field solutions in Eqs. (3) and (4) at $x=\tfrac{d}{2}^{+}$ (on the side of the air gap) to derive a surface impedance that represents the effective boundary condition for the WM interface,
@@
This impedance vanishes for $g=0$, corresponding to a perfect electric conductor, and approaches infinity as $g$ tends to $\lambda/4$ matching the wave impedance of a perfect magnetic conductor.
II.2 Fundamental mode of WM resonator
In this section, we investigate the TM modes of a 1D WM resonator illustrated by Fig. 2 for the limit cases of $g=0$ and $g=\lambda/4$. We assume no field variation along the $z$-axis for the fundamental WM resonator mode (i.e. $k_{z}=0$) and consider a square resonator cross-section ( implying $k_{x}=k_{y}=k_{t1}$). Using these assumptions, we express the electromagnetic field within the effective WM as
@@ II.2 Fundamental mode of WM resonator
For the fundamental TM mode within the effective WM region, defined by $|x|,\>|y|\leq\tfrac{d}{2}$, we define the wave number as
@@
while we define the wave number in the air gap region, for instance where $x>\tfrac{d}{2},|y|\leq\tfrac{d}{2}$, as
@@
It is important to note that in this model, the WM resonator has no corners, which may have impact for the real resonator. Despite this approximation, we can employ the one-dimensional solution to estimate the two-dimensional behavior, analogous to Marcatili’s method for dielectric rectangular waveguides [26]. This approximation is valid when the field is primarily concentrated in the central region.
@@
This dispersion equation yields the resonance frequencies of the TM modes. Notably, using Eq. (9), we can determine the resonance frequency of the fundamental TM mode. For the fundamental mode, when $g=\lambda/4$, the second term in Eq. (10) vanishes, implying that $k_{t1}=0$. Consequently, $\epsilon=0$ and $k_{0}=k_{p}$, leading to a constant electric field according to Eq. (6). It is worth mentioning that Eq. (10) is similar to the dispersion equation derived in Ref. [16]. However, while Ref. [16] achieves field homogeneity in a single direction by using a central metallic waveguide operating at its cutoff frequency, our approach requires a WM in the central region of the cavity operating at the plasma frequency to ensure field homogeneity in both transverse directions.
@@ II.3 Results
Figure 3: Electric field distribution in a wire medium (WM) resonator: (a) Regular WM resonator with no air gap ($g=0$), and (b) WM resonator with an air gap of thickness $g=\lambda_{p}/4$. The introduction of the air gap ($g=\lambda_{p}/4$) results in a more uniform electric field distribution.
@@
Figure 4: Wire medium (WM) resonator electric field profile for different air gap thicknesses, $g$. The WM extends from $x=-\tfrac{d}{2}$ to $x=\tfrac{d}{2}$. A uniform electric field in the WM region is observed at $g=\lambda_{p}/4$.
@@
Figure 5: Characteristics of wire medium (WM) resonator fundamental TM mode. a) Resonance frequency and b) form factor for different air gap thicknesses, $g$. A form factor of approximately 0.92 is observed at $g=\lambda_{p}/4$ for an effective WM (theory and numerics) while a form factor of 0.89 is observed for the physical WM (numerics). Point P corresponds to the air gap thickness used in the experimental section.
@@
We used full-wave simulations in a COMSOL Multiphysics eigenmode solver to validate the applicability of our 1D theory to a 2D cavity. In both our analytical calculations and simulations, the cavity comprised an array of 10$\times$10 infinitely long wires with wire radius $r_{0}$ = 0.5 mm and inter-wire spacing $a$ = 10 mm. We adopted the generalized quasi-static approach from Ref. [27] to determine the requisite plasma frequency for the theoretical model.
@@
Figure 3 illustrates the arbitrary normalized electric field distribution for the fundamental TM mode over the transverse ($xy$) plane of the cavity for two specific cases: $g=0$ and $g=\lambda_{p}/4$. Our numerical and analytical findings confirm that introducing a $\lambda_{p}/4$ air gap homogenizes the electric field distribution in a finite-cross section sample of the WM.
@@
The electric field profile along the $x$-axis for the fundamental TM mode of a WM resonator varies with the gap width, $g$, as shown in Fig. 4. The electric field homogeneity improves with increasing $g$ until reaching a maximum at $g=\lambda_{p}/4$. Introducing a quarter-wavelength air gap between the effective wire medium and the PEC walls significantly flattens the electric field profile, confining the field decay primarily to the newly added $\lambda_{p}/4$ region. This coincides with a pronounced magnetic field enhancement within the same gap, offering potential for readout applications. For the gap exceeding $\lambda_{p}/4$, the field concentrates near the resonator walls, and the underlying assumption of field confinement within the wire medium becomes not adequate.
@@
Figure 5 illustrates the resonant frequency (Fig. 5a) and form factor (Fig. 5b) of the fundamental mode in such WM resonator as a function of the air gap, $g$. The resonant frequency remains nearly constant for $40\leq g\leq\lambda_{p}/$ 4, aligning with the plasma frequency at $g=\lambda_{p}/4$. Conversely, the form factor progressively increases, reaching a maximum at an air gap of approximately $g=\lambda_{p}/4$. While the resonant frequency exhibits excellent agreement between analytical and numerical results, coming from both a physical WM and an effective WM, within this range, a slight discrepancy arises in the form factor. This error results from the evident difference between the effective and physical (real) WM. In the effective-medium model there are no wires and the field does not vanish inside this medium, whereas in the physical WM it vanishes within every wire – see the dark spots in the bottom panel of Fig. 3b. In accordance with numerical calculations of the mode field within a single unit cell of our WM the maximally achievable form factor corresponding at 8.9 GHz to $g=\lambda_{p}/4=$8.4 mm is equal to 0.92, as well as in the analytical model. However, the simulation of the physical WM corresponding to the bottom panel of Fig. 3b shows the form factor 0.89. Despite this, the maximum form factor occurs at nearly the same air gap value in both numerical simulations and the analytical model. For smaller gaps the agreement between the analytical and numerical results corresponding to physical WM is good, and is excellent for the numerical simulations of the effective WM sample surrounded by the PEC walls. Thus, these results fully justify the model of an effective WM sample for $g\leq\lambda_{p}/4$. For $g>\lambda_{p}/4$, the agreement between the analytical model and numerical simulations disappears due to the effect of the field concentration in the air gap. This effect invalidates the initial assumption of field confinement within the WM sample. However, the agreement between the results obtained for the effective WM sample and for the array of separate wires keeps.
@@
It is evident that the maximal form factor corresponds to the perfect magnetic wall effectively located on the WM-air interface. The PEC wall shifts the phase of a reflected wave by $\pi$, the air gap grants an extra $\pi$, so the phase of the reflected wave does not change. Therefore, the mode inside the WM sample is the standing wave shaped by four perfect magnetic boundaries surrounding the sample. Note that this effective boundary condition can be achieved through various methods that introduce an extra phase shift of $\pi$ between incident and reflected waves, as detailed in [16, 17, 18]. This flexibility may help the TM000 modes implementation in innovative plasma haloscope designs that require tunability.
III Experiment
@@
Figure 6: Prototype WM resonators with different wall spacing. The resonators consist of 10 x 10 arrays of copper wires with a period of $a=1$ cm. The first resonator is a regular resonator with a distance of $a/2$ between the wires at the edge and the walls, and the second one is $\lambda/4.2$-shifted with a distance of approximately $a/2+\lambda/4.2$ between the wires at the edge and the walls. (a) Unfinished state of the regular resonator. Copper wires (radius 0.5 mm) are inserted into the FR-4 PCB walls with a period of 1 cm. (b) Finished prototype being scanned through pre-drilled holes by a movable antenna.
@@
We fabricated two prototypes to validate the claimed effect. The first prototype represents the regular (conventional) configuration, see for example Ref. [14], where the air gap between the array of wires and the cavity walls is equal $a/2$. The second prototype incorporates a configuration we specially optimized for the cubic resonator to enhance field homogeneity and to avoid situations where the air gap exceeds $\lambda_{p}/4$, as this would lead the field to localize mainly near the cavity walls (see Fig. 4). In this 3D case, the designed gap is slightly smaller than the quarter-wavelength. So, for the optimized resonator, the edge wires of the array are distanced from the walls by ($a/2+\lambda_{p}/4.2$). For the 2D resonator, this configuration would correspond to point P in Fig. 5. Below we call our optimized cubic resonator the $\lambda_{p}/4.2$-shifted one.
@@
We fabricated the prototype enclosures using single-sided metal FR-4 PCBs of $1$ mm thickness. We arranged these PCBs into cubes with dimensions of 10 cm x 10 cm x 10 cm for the regular resonator and 11.6 cm x 11.6 cm x 11.6 cm for the $\lambda/4.2$-shifted resonator. To complete the resonators, we inserted and welded one hundred copper wires, each with a radius of 0.5 mm, into the PCB cubes with a periodic spacing of $a=1$ cm, as shown in Fig. 6a.
@@
Figure 7: Experimentally measured $\mathrm{S}_{21}$ parameter for wire media resonators: (a) Regular resonator and (b) $\lambda_{p}/4.2$-shifted resonator. Measurements were performed by placing the movable monopole antenna in the central hole. The resonance frequency of the fundamental TM mode in the regular resonator ($f_{110}$) is $9.11$ GHz (experiment) and $9.10$ GHz (simulation), while in the $\lambda_{p}/4.2$-shifted resonator ($f_{000}$) it is $8.943$ GHz (experiment and simulation).
We measured the S-parameters using a Rohde &\&& Schwarz ZVB20 vector network analyzer (VNA) across a frequency range of 8.75-9.75 GHz. The fixed antenna, located on the bottom plate of each prototype, was connected to one port of the VNA, while the movable antenna on the top plate was connected to the other port. To measure the $\mathrm{S_{21}}$, we positioned the monopole antenna in the array of holes. Fig. 7 displays the measured $\mathrm{S_{21}}$ when the movable antenna was placed in the central hole of each prototype. We observe that the spectrum of the $\lambda_{p}/4.2$-shifted resonator is red-shifted relative to that of the regular resonator, as predicted by theory for the fundamental mode. Meanwhile, the amplitude levels and shape of the spectrum change slightly, indicating that the quality factor of the resonator is not significantly affected when an air gap of $\lambda_{p}/4.2$ is introduced.
We conducted full-wave simulations of cubic resonators, corresponding to the experimental prototypes, using CST Studio. Copper was modeled with a conductivity of $\sigma=5.7\times 10^{7}\;\mathrm{S/m}$, while the FR-4 layers of the PCBs were represented as lossy dielectrics with a dielectric constant of $\varepsilon=4.3$ and a loss tangent of $\tan{\delta}=0.025$. We measured the $\mathrm{S_{21}}$ parameter and positioned a field monitor at the first resonance frequency for each resonator.
@@
We characterized the operation of the two resonators under comparison by examining the first peaks of the $\mathrm{S_{21}}$-parameter spectra. These peaks correspond to the $\mathrm{TM_{110}}$ mode for the regular resonator and the quasi-$\mathrm{TM_{000}}$ mode for the $\lambda_{p}/4.2$-shifted resonator, as shown in Fig. 7. We do not consider high-order modes, as they are not relevant for our purpose of achieving the highest field homogeneity (see e.g. in Refs. [17, 14]). Fig. 8 presents the measured electric field distributions for both prototypes at the corresponding frequencies ($f_{000}$ for the $\lambda_{p}/4.2$-shifted resonator and $f_{110}$ for the regular one). We observe a significant improvement in field homogeneity for the $\lambda_{p}/4.2$-shifted resonator compared to the regular resonator. This observation is further supported by the measured form factors, with a value of $0.77$ (experiment), $0.80$ (simulation) for the $\lambda_{p}/4.2$-shifted resonator compared to $0.61$ (experiment), $0.64$ (simulation) for the regular resonator.